# Remote Sensing Image Intelligent Interpretation with the Language-Centered Perspective: Principles, Methods and Challenges

### v1

@@ II-B The Potential of Applying LLMs as the Global Workspace
Figure 4: Paradigm Shift in Remote Sensing Image Interpretation: From Vision-Centered to Language-Centered
II-C Redefining Intelligent Interpretation of Remote Sensing Imagery
Currently, the mainstream approach to intelligent interpretation of remote sensing imagery follows a vision-centered passive perception paradigm, typically employing an end-to-end inference mechanism: input data is processed by a model, which then produces interpretation results based on pre-trained parameters. Under this paradigm, the model is usually tightly coupled with the specific task, meaning that an independent model must be trained for each individual task, and that model relies heavily on data from a particular application scenario. This highly specialized design significantly limits the model’s generalization ability. Especially when dealing with complex tasks involving multiple decision-making processes, it often becomes necessary to design and train separate models for each decision flow to meet varying task requirements. As a result, the overall system construction cost increases substantially, and the flexibility to adapt to changing scenarios is greatly reduced.
In contrast, bottom-up (perception-driven cognition) and top-down (cognition-guided perception) processes represent two fundamental pathways through which humans understand the world. Combining these two paths enables the formation of a ”Perception + Cognition” cyclic framework. This framework not only allows for the in-depth extraction of visual information at different levels of granularity, but also supports the construction of a multi-task reasoning chain by associating perceptual input with existing knowledge, thereby enabling complex visual reasoning tasks.
As illustrated in Fig.4, traditional ship target recognition methods rely on a single visual model and follow an end-to-end production pipeline, where task forms are fixed. In contrast, a language-centered approach to intelligent interpretation of remote sensing imagery allows users to interact with the model through natural language and actively assess the progress of interpretation tasks. When key information is missing, the model can interact with sensors to acquire additional perceptual data, thereby significantly enhancing the flexibility and adaptability of the interpretation system.
Therefore, this paper redefines the process of intelligent interpretation of remote sensing imagery as follows: it is an interpretation process centered on language-organized interaction among tasks, sensors, and models, and driven by cyclic, multi-round dialog-based reasoning. This process integrates perception and cognition to support flexible task configurations. Remote sensing image interpretation should no longer be centered on vision-driven passive perception, but should instead be realized through a multi-level framework that combines perception and cognition.
This framework effectively integrates perception-driven visual processing with cognition-driven semantic reasoning, enabling the interpretation system to go beyond basic image analysis and leverage external knowledge for task reasoning and information completion. Such a ”perception + cognition” interpretation framework not only has the potential to enhance the generalization capabilities of interpretation models but also provides an effective solution for addressing complex interpretation tasks.
@@ II-D Advantages of Language-Centered Intelligent Interpretation of Remote Sensing Imagery
The language-centered intelligent interpretation paradigm emulates the fundamental way in which humans understand the world—by integrating the perception space, knowledge space, task space, and action space into a cyclic interpretation model capable of effective interaction with the physical world. In this paradigm, language is not merely a tool for information transmission, but serves as the central hub for knowledge integration, coordination, and decision-making. Through a language-driven cyclic interpretation process, the model can handle complex interpretation tasks in open environments, enabling a shift from passive environmental perception to active environmental interaction in remote sensing image interpretation.
The rapid advancement of LLMs makes this paradigm increasingly feasible. With their abilities in unified representation, knowledge association, and reasoning and decision-making, LLMs overcome the inherent limitations of vision-based models—limitations stemming from the tightly coupled nature of data, models, and tasks in traditional training pipelines. As a result, language-centered models demonstrate the following advantages in intelligent remote sensing image interpretation:
@@
Language-centered interpretation models offer advantages in integrating multimodal perceptual data
Due to differences in their intrinsic mechanisms (e.g., spatial resolution, spectral characteristics) and extrinsic factors (e.g., temporal and spatial coverage), different modalities are often highly heterogeneous at the structural level, posing significant challenges for alignment and collaborative interpretation. Numerous studies have found that in human cognition, information from various modalities—such as auditory, tactile, olfactory, and visual—is ultimately conceptualized and represented through language. Language serves not only as a carrier of perception and understanding but also as the medium for association, reasoning, and expression. This suggests that language not only encodes our perception of the world with precision but also implicitly establishes associations across different modalities. Although remote sensing data types such as optical, infrared, hyperspectral, SAR, and LiDAR differ in both intrinsic mechanisms and acquisition conditions, they can all be considered semantically as different views of the same geo-object. Recent multimodal language model research—such as CLIP[7], LanguageBind[21], and AllSpark[22] has validated the feasibility of embedding diverse modalities into a shared language space. Therefore, language-centered interpretation models possess the potential and advantages to align multimodal perceptual data semantically via the language modality and to effectively fuse heterogeneous information across modalities.
Language-centered interpretation models possess the ability to learn in open environments, enabling them to quickly adapt and integrate new information when encountering unfamiliar scenes or data
In contrast, traditional vision-based models primarily rely on supervised learning through the mapping between massive amounts of remote sensing imagery and a limited set of semantic annotations. This approach suffers from inherent theoretical limitations, as a finite set of semantic labels cannot provide sufficient information to thoroughly describe geo-objects and their relationships—thus restricting the model’s understanding of the world to the boundaries defined by the labeled knowledge. As a result, remote sensing vision models often struggle to perform effectively when faced with unlabeled or novel scenarios, making it difficult for them to generalize beyond predefined categories. In stark contrast, LLMs encode human knowledge that is inherently expressed in language. By associating perceptual information with a vast and structured knowledge base, LLMs can overcome the constraints imposed by limited annotations and significantly expand the range of learnable knowledge. For example, in ship recognition tasks, even when encountering previously unseen ship categories, a language-centered model can identify key component features from the image and reason about the specific type of vessel by associating those features with relevant prior knowledge. Specifically, different classes of ships, such as aircraft carriers, typically exhibit stable and class-distinguishing features. For instance, the number of ship islands can be used to distinguish the Queen Elizabeth-class aircraft carrier from other carriers. This is made possible because language enables compositional reasoning, allowing the model to associate various component features and derive logical conclusions.
@@
Language-centered interpretation models offer clear advantages in handling complex interpretation tasks
Leveraging the conceptual relationships encoded within LLMs, such models can decompose complex tasks into simpler subtasks and perform reasoning and decision-making by integrating and associating diverse information within the language space. In contrast, vision-centered interpretation models typically require a separate, dedicated model for each complex task. This not only increases the complexity of model development and maintenance but also limits each model’s applicability to a narrow task scope, making it difficult to adapt to rapidly changing interpretation demands. For example, in remote sensing monitoring of flood disasters, an LLM can first analyze low-resolution remote sensing images collected by various sensors to identify the affected flood zones. It can then match these zones with spatial distribution maps of different functional areas—such as agricultural or industrial zones—to infer the degree of damage in each area. Based on this preliminary damage assessment, the LLM can dynamically mobilize relevant sensing resources, such as high-resolution imagery sensors, meteorological sensors, and geological sensors, to acquire more detailed and real-time remote sensing data for further analysis of the flood’s impact on each region. In parallel, by combining historical data with real-time sensor inputs, the model can assess the potential risk of secondary disasters, such as landslides or mudflows. Through this cyclic, step-by-step reasoning and decision-making process, the model can generate comprehensive predictions of both the flood’s impact and the likelihood of associated secondary hazards.
Language-centered interpretation models have the advantage of actively acquiring perceptual information
By deeply understanding the interpretation task, these models can assess whether the available perceptual information is sufficient and proactively seek additional data when needed. This capability stems from the model’s mechanisms for knowledge association and reasoning-based decision-making, allowing it to evaluate the relevance and completeness of information during interpretation and expand its knowledge scope when necessary to meet the demands of complex tasks. In contrast, traditional vision-based models are limited to the perceptual information contained within the imagery itself. When such information is insufficient or when complex scenarios require external knowledge support, vision models often struggle to adapt. For instance, in the task of observing and analyzing ship types, conventional imagery may fail to provide sufficient recognizable features due to weather conditions (e.g., clouds or fog) or suboptimal observation angles. In such cases, a language model can analyze the perceptual input—such as blurred images or missing hull numbers—and reason about the next best observation action to fulfill the interpretation goal. The model can then guide the system to perform more targeted sensing operations, such as issuing a command for a drone to switch to a high-resolution optical camera and move closer to the target for clearer imaging. Through this process, the language model can actively acquire critical information, enabling more effective completion of the interpretation task.
@@
In summary, compared to traditional vision-centered interpretation models, language-centered models demonstrate clear advantages. Leveraging their rich knowledge association capabilities and powerful reasoning abilities, language models can actively acquire and integrate multi-source data, effectively addressing the challenges posed by open remote sensing scenarios and complex interpretation tasks. The shift from a vision-centered to a language-centered paradigm broadens the scope of intelligent remote sensing image interpretation and enhances the model’s adaptability and interpretive accuracy in contexts characterized by multi-source data, open environments, and complex task demands.
@@ III A Language-Centered Framework for Intelligent Interpretation of Remote Sensing Imagery
Although the language-centered framework for intelligent interpretation of remote sensing imagery demonstrates significant advantages over traditional vision-centered approaches—particularly in handling complex and open remote sensing scenarios with greater flexibility, interactivity, and reasoning capability—it also faces a series of challenges in practical implementation, as illustrated in Fig.5. First, remote sensing imagery in the perception space encompasses multiple modalities, and the differences in spatial and temporal resolution, coverage, and spectral characteristics across these modalities make it difficult to unify their representations within a global workspace. Second, the perceptual information from the perception space, the interpretation tasks from the task space, and the internal and external prior knowledge in the knowledge space are often difficult to effectively associate. Finally, the knowledge elements themselves often exhibit complex logical relationships, making it challenging to perform accurate reasoning and to generate reliable decisions for subsequent actions.
In summary, intelligent interpretation of remote sensing imagery faces significant challenges in three key areas: multimodal unification, knowledge association, and reasoning and decision-making. Multimodal unification concerns the harmonization of data representations across different modalities and the alignment of their semantic meanings. The core of knowledge association lies in accurately understanding the interpretation task and retrieving relevant knowledge with precision. Reasoning and decision-making focus on dynamically analyzing interpretation progress based on perceptual inputs and associated knowledge, and on effectively planning observation strategies to optimize the interpretation process. This section will first provide an in-depth analysis of these core challenges, followed by a review of the current research status and recent advancements in each of these subdomains.
@@ III-A1 Unified Multimodal Representation
Non-linguistic modalities differ significantly from language modalities in terms of data representation formats, making it difficult to embed them into a shared semantic space. For example, RGB, MSI, and SAR imagery are typically represented as dense 2D matrices, while LiDAR point cloud data is stored as sparse 3D point sets. These structural inconsistencies can lead to loss of perceptual information when non-linguistic modalities are mapped into a 1D language-based semantic space, thereby affecting the effectiveness of multimodal fusion. Therefore, one of the fundamental challenges for language-centered unified representation is how to standardize the representation of non-linguistic modalities in a way that preserves their perceptual content while aligning them with language-based representations.
@@
Language modalities carry human abstract conceptual systems, while non-language modalities are direct samplings of the physical world. This fundamental difference means that aligning their semantics typically requires additional human prior knowledge. However, in the field of remote sensing, non-language modality data is vast and rich in spatiotemporal information, making it extremely costly and labor-intensive to build precise alignment between language and non-language modalities through manual annotation. Furthermore, there are often significant mismatches in granularity and informational content between the two modalities. For instance, the term ”ship” in natural language typically refers to the entire vessel, whereas in remote sensing imagery, the representation of a ship may consist of various individual visual components. Direct cross-modal alignment without proper abstraction may result in semantic ambiguity. Thus, a core challenge for language-centered multimodal fusion is how to establish accurate semantic alignment mechanisms while minimizing reliance on manual annotation.
III-A2 Knowledge Association
The global workspace contains a complex composition of knowledge, and the ability to associate and integrate knowledge from different sources plays a critical role in solving complex remote sensing image interpretation tasks. Knowledge association is essential for enabling the system to make correct reasoning and decisions. Therefore, achieving accurate knowledge association is one of the primary challenges in building a language-centered intelligent interpretation system for remote sensing imagery. Knowledge association is a complex process that requires establishing effective links among perceptual information, task requirements, and both internal and external knowledge sources. Specifically, it involves two key steps: task understanding and knowledge retrieval.
@@ III-A2 Knowledge Association
Task understanding is the prerequisite for accurate knowledge association. A remote sensing interpretation system must deeply comprehend the task requirements and relate them to perceptual information to determine which pieces of information are most relevant. For example, in a fine-grained ship classification task, different types of ships within the same general category often differ only in specific subtle features. Therefore, the interpretation system must leverage internal knowledge to first identify the coarse category of the target ship, and then analyze the image content to determine which visual features—such as the position or number of ship islands, or the hull number—are most helpful for fine-grained classification. This enables the system to more efficiently guide the execution of subsequent recognition steps.
@@
Knowledge retrieval is key to accurately extracting relevant information. Remote sensing interpretation systems must not only utilize the knowledge encoded within the model itself but also possess the ability to extract task-relevant knowledge from external multimodal sources. Continuing with the example of fine-grained ship classification, once the system has fully understood the task requirements, it must retrieve and extract information related to ship subcategories and their distinctive features from both internal and external knowledge sources. For instance, the Queen Elizabeth-class aircraft carrier is equipped with two ship islands—this type of feature can serve as crucial evidence in subsequent reasoning and decision-making processes and directly impacts the system’s interpretation performance. In practical terms, knowledge retrieval involves two core technical challenges: efficient knowledge encoding and reliable knowledge retrieval. The former requires the model to embed various forms of data—such as structured data, unstructured text, or temporal information—into its internal parameters, while ensuring that the information is well-organized and meaningfully associated to avoid knowledge confusion. The latter demands that the model be able to rapidly locate and retrieve information highly relevant to the current task from a vast internal and external knowledge space, by leveraging the complex associative relationships among knowledge elements.
III-A3 Reasoning and Decision-Making
Accurate and efficient reasoning and decision-making are key to driving the cyclic interactive interpretation process. Within the global workspace, reasoning and decision-making require the interpretation system to integrate task-relevant internal knowledge during the perception process. The system must first assess whether the currently available information is sufficient to accomplish the interpretation task, and then identify any missing information and formulate an optimal next-step sensing strategy. Through this process, the interpretation system can proactively select regions or targets that require further observation, thereby improving both sensing efficiency and interpretation accuracy.
Assessing the gap between external information and task objectives
One of the primary challenges faced by interpretation systems is how to evaluate the gap between the available external information and the requirements of the task. This gap assessment largely determines the system’s cognitive capability and perceptual efficiency. Only when the system can accurately assess the relevance and adequacy of the current external information can it make reasonable decisions for subsequent planning. The process of evaluating this gap is essentially a multi-level and complex computation. It requires both a precise understanding of the task objectives and a comprehensive analysis of the characteristics and variations in the external data. If this evaluation is inaccurate, the system’s reasoning path may deviate from the correct decision-making direction, leading to biased interpretation results and ultimately compromising task performance. Therefore, designing an efficient and accurate mechanism for information gap assessment—one that enables the system to maintain stable and precise judgment in complex and dynamic information environments—is a crucial prerequisite for achieving active perception.
Reasoning path optimization and action decision-making
A central challenge lies in designing an efficient decision-making mechanism that ensures the reasoning path leads to optimal perceptual efficiency. The purpose of the reasoning path is to progressively narrow the gap between external information and task objectives, allowing the system to dynamically adapt to environmental changes as it moves closer to its goal. However, in real-world physical environments, external conditions are often highly dynamic and influenced by numerous factors—such as the movement of target objects, changes in lighting, or variations in weather conditions within the sensing range. These variations cause perceptual information to continuously change in both space and time. In such contexts, the interpretation system cannot rely on fixed, pre-defined task strategies for information acquisition. Therefore, the system must possess flexible, adaptive capabilities that allow it to adjust its reasoning path in real time based on environmental changes, ensuring that it can still capture critical information effectively in dynamic conditions. As such, designing an adaptive decision-making mechanism that meets the practical demands of real-world tasks and guides the system to dynamically plan its reasoning and action path is another core research issue for achieving efficient reasoning and decision-making.

### v2

@@ II-B The Potential of Applying LLMs as the Global Workspace
Figure 4: Paradigm Shift in Remote Sensing Image Interpretation: From Visual-Centered to Language-Centered
II-C The Paradigm Shift in Remote Sensing Interpretation: From Visual-Centered to Language-Centered
Currently, the mainstream approach to remote sensing intelligent interpretation remains dominated by a visual-centered passive perception paradigm, which executes tasks through end-to-end inference mechanisms. Although recent visual-based geospatial foundation models (GFMs)[16, 48, 177] have significantly advanced feature extraction and heterogeneous alignment across multimodal data, they remain constrained by three inherent limitations as task complexity and reasoning depth increase:
  • Semantic Gap: These models rely on data alignment rooted in physical consistency, making it difficult to bridge the semantic divide. Recent works such as DOFA[182] and msGFM[45] essentially rely on spatiotemporal consistency to construct feature associations. Although effective at integrating heterogeneous visual signals such as optical, SAR, hyperspectral, and infrared imagery, this mechanism struggles to bridge the gap between ’pixel-level features’ and ’abstract concepts.’ Particularly when addressing abstract semantics that lack direct physical manifestations (e.g., ’factory abandonment due to economic recession’ or ’complex tactical intents’), purely visual models are unable to transcend the cognitive barrier from low-level signals to high-level logic, primarily because of the lack of a symbolic semantic interface.
  • Closed Knowledge System: It is challenging to incorporate open-ended external knowledge. The knowledge within the VFM is derived solely from the pixel patterns in the training data. This “what you see is what you get” learning style constructs a closed system, preventing the direct utilization of vast external knowledge bases (e.g., historical observations, geographic laws, or expert expertise). In contrast, language models possess an innate ability to connect with world knowledge. Lacking such symbolic knowledge injection channels, visual models often show a deficit in generalization when handling complex tasks requiring background reasoning, such as long-tail object recognition or disaster causality analysis.
  • Passive Perception: Limited by a static mapping paradigm. Regardless of their feature extraction strength, existing visual models exhibit unidirectional “input image to output result” mapping. They lack agency; they cannot understand complex, unstructured intents such as an agent, nor can they perform reflection or proactive planning (such as invoking external tools or adjusting observation perspectives) when information is insufficient. This lack of interactivity renders visual-centered models passive data analysis tools rather than intelligent cognitive systems capable of autonomous problem solving.
Consequently, the interpretation process of the visual-centered paradigm (as illustrated in Fig. 4(a) is restricted to judgments based on local, explicit clues within the current field of view. It lacks the capacity for cross-scale association, intent understanding, and information completion. When visible information is insufficient, the model cannot infer strategic actions such as “further observation required” or “larger-scale imagery needed,” thus leading to erroneous results. As shown in the figure, owing to missing visual cues, the model misidentifies “water” in the lower-left corner as “farmland.”
In contrast to passive perception, the language-centered interpretation paradigm emphasizes the organic integration of bottom-up perception and top-down semantic reasoning (Fig. 4(b)). In this framework, perception provides explicit clues, while language-led cognition infers results based on semantic structure and directs further observation. As illustrated, the model initially struggles to classify the lower-left area. The language reasoning module then proactively requests “more perceptual information,” prompting the system to retrieve an enlarged patch from the database. This additional visual context enables the model to accurately identify the true category.
In this language-centered paradigm, perception-driven visual processing and cognition-driven semantic reasoning are effectively integrated. This integration allows the system to move beyond basic image analysis to perform complex reasoning and information completion using language-organized knowledge. This “perception (extracting image information), cognition (water or farmland?), decision (finding more information) and action (searching the knowledge base)” mechanism not only enhances generalization but also enables the system to robustly handle interpretation tasks with high structural complexity and rich semantics.
@@ II-D Advantages of Language-Centered Intelligent Interpretation of Remote Sensing Imagery
The language-centered intelligent interpretation paradigm mimics the fundamental human way of understanding the world. By integrating the spaces of perception, knowledge, task, and execution, it constructs a cyclic interpretation mode capable of effective interaction with the physical world. In this mode, language transcends its role as a mere communication tool, functioning instead as the central hub for knowledge integration, orchestration, and decision-making. Through this language-driven cyclic process, the model can address complex tasks in open environments, achieving transformation from passive environmental perception to active environmental interaction. The rapid advancement of LLMs has rendered this paradigm feasible. The capabilities of LLMs in unified representation, knowledge association, and reasoning-based decision-making overcome the inherent limitations of traditional visual models caused by the strong coupling of data, architectures, and tasks. Consequently, language-centered models demonstrate the following advantages in remote sensing interpretation:
@@
Language-centered interpretation models demonstrate significant advantages in integrating multimodal perceptual data
Stemming from variations in intrinsic imaging mechanisms (e.g., spatial resolution and spectral characteristics) and extrinsic acquisition conditions (e.g., spatiotemporal coverage), multimodal data exhibit high heterogeneity at the structural level. The traditional visual-centered paradigm struggles to bridge this physical gap, often necessitating strict pixel-level spatiotemporal alignment to achieve effective fusion. In contrast, the language-centered paradigm effectively circumvents this bottleneck. Although remote sensing data, ranging from optical and infrared to SAR and LiDAR, diverge sharply in terms of physical attributes, from a language-centered perspective, they are essentially different representations of the same geo-object. Consequently, the language modality serves as a unified semantic anchor, enabling deep association and unified representation of multimodal data at the semantic level without the need for strict physical alignment.
Language-centered interpretation models possess superior open-environment learning capabilities, enabling flexible adaptation to unknown scenes and the rapid integration of new knowledge
In contrast, traditional visual models rely primarily on the mapping between “massive imagery and limited labels” as their supervision signal. This paradigm suffers from fundamental theoretical limitations: discrete and finite semantic labels fail to encapsulate the rich information regarding ground objects and their spatiotemporal relationships. This shortcoming inevitably confines the model’s cognitive boundaries within a predefined closed-set knowledge domain, resulting in severely restricted generalization capabilities when identifying unlabeled or novel scenes. Conversely, large language models (LLMs) associate visual perceptual information with expansive open-world knowledge systems. This feature allows models to effectively overcome the cognitive bottlenecks imposed by limited annotation categories, achieving zero-shot understanding of unseen concepts.
@@
Language-centered interpretation models demonstrate significant advantages in handling complex cognitive tasks
The traditional visual-centered paradigm typically adopts a “one-model-per-task” architecture, necessitating the construction of independent, specialized models for distinct tasks such as object detection and change detection. This approach not only incurs high development and maintenance costs but also creates severe semantic barriers between models, making it difficult to address complex interpretation requirements that demand multistep reasoning. In contrast, language-centered models transcend these rigid task boundaries. Leveraging the powerful logical reasoning capabilities and chain-of-thought (CoT)[176] mechanisms of LLMs, these models automatically decompose complex macrointerpretation tasks (e.g., “analyzing port traffic”) into executable atomic sub-tasks (e.g., “identifying vessels,” “counting objects,” and “assessing status”). By mapping multimodal perceptual information into a unified linguistic space, the model can perform logical deduction and decision-making akin to the capabilities of human experts. This paradigm shift from “task-specific mapping” to “general-purpose logical planning” not only drastically reduces the complexity of multitask systems but also endows interpretation models with “generalist” capabilities to flexibly handle diverse, nonstandard requirements in rapidly changing open environments.
@@
In summary, compared with traditional visual-centered interpretation models, language-centered models demonstrate clear advantages. Leveraging their rich knowledge association capabilities and powerful reasoning abilities, language models can actively acquire and integrate multisource data, effectively addressing the challenges posed by open remote sensing scenarios and complex interpretation tasks. The shift from a visual-centered to a language-centered paradigm broadens the scope of intelligent remote sensing image interpretation and enhances the model’s adaptability and interpretive accuracy in contexts characterized by multisource data, open environments, and complex task demands.
@@ III A Language-Centered Framework for Intelligent Interpretation of Remote Sensing Imagery
Drawing upon GWT, this section proposes a language-centered intelligent interpretation paradigm for remote sensing imagery, aiming to construct a complete interpretation closed loop spanning from multimodal perception to high-level cognition and, finally, to action feedback (as illustrated in Fig. 5). This paradigm abstracts complex interpretation tasks into three progressive cognitive stages. First, heterogeneous multimodal perceptual information ($P$), such as optical and radar data, is mapped onto a unified global workspace through an alignment mechanism to establish a unified perceptual representation. Second, dynamic associations among perceptual information, remote sensing expert knowledge ($K$), and task requirements ($T$) are facilitated within the workspace to conduct deep semantic reasoning. Finally, decision-making is performed based on reasoning results and translated into specific executable actions ($A$) to advance the interpretation process.
However, while this paradigm offers a clear logical framework for the synergy of “perception-cognition-action”, its practical implementation faces critical challenges from physical alignment and cognitive reasoning to decision execution. Specifically, these challenges include overcoming the physical heterogeneity of multimodal data to achieve a unified representation; ensuring precise knowledge association and reasoning within complex knowledge bases; and selecting optimal solutions from ambiguous inferences for efficient decision-making and execution. In this section, these three core issues are explored in depth.
@@ III-A1 Unified Multimodal Representation
Nonlinguistic modalities differ significantly from language modalities in terms of data representation formats, making it difficult to embed them into a shared semantic space. For example, RGB, MSI, and SAR imagery are typically represented as dense 2D matrices, while LiDAR point cloud data are stored as sparse 3D point sets. These structural inconsistencies can lead to a loss of perceptual information when nonlinguistic modalities are mapped onto a 1D language-based semantic space, thereby affecting the effectiveness of multimodal fusion. Therefore, one of the fundamental challenges for language-centered unified representation is how to standardize the representation of nonlinguistic modalities in a way that preserves their perceptual content while aligning them with language-based representations.
@@
Language modalities carry human abstract conceptual systems, whereas nonlanguage modalities are direct samplings of the physical world. This fundamental difference means that aligning their semantics typically requires additional human prior knowledge. However, in the field of remote sensing, nonlanguage modality data are vast and rich in spatiotemporal information, making it extremely costly and labor intensive to achieve precise alignment between language and nonlanguage modalities through manual annotation. Furthermore, there are often significant mismatches in granularity and informational content between the two modalities. For instance, the term “ship” in natural language typically refers to the entire vessel, whereas in remote sensing imagery, the representation of a ship may consist of various individual visual components. Direct cross-modal alignment without proper abstraction may result in semantic ambiguity. Thus, a core challenge for language-centered multimodal fusion is how to establish accurate semantic alignment mechanisms while minimizing reliance on manual annotation.
III-A2 Knowledge Association and Reasoning
The global workspace contains a complex composition of knowledge, and the ability to associate and integrate knowledge from different sources plays a critical role in solving complex remote sensing image interpretation tasks. Knowledge association is essential for enabling the system to make correct reasoning and decisions. Therefore, achieving accurate knowledge association is among the primary challenges in building a language-centered intelligent interpretation system for remote sensing imagery. Knowledge association is a complex process that requires the establishment of effective links among perceptual information, task requirements, and both internal and external knowledge sources. Specifically, it involves two key steps: task understanding and knowledge retrieval.
@@ III-A2 Knowledge Association and Reasoning
Task understanding is a prerequisite for accurate knowledge association. A remote sensing interpretation system must deeply comprehend the task requirements and relate them to perceptual information to determine which pieces of information are most relevant. For example, in a fine-grained ship classification task, different types of ships within the same general category often differ only in specific subtle features. Therefore, the interpretation system must leverage internal knowledge to first identify the coarse category of the target ship and then analyze the image content to determine which visual features–such as the position or number of ship islands or the hull number–are most helpful for fine-grained classification. This approach enables the system to more efficiently guide the execution of subsequent recognition steps.
@@
Knowledge retrieval is key to accurately extracting relevant information. Remote sensing interpretation systems must not only utilize the knowledge encoded within the model but also possess the ability to extract task-relevant knowledge from external multimodal sources. Continuing with the example of fine-grained ship classification, once the system has fully understood the task requirements, it must retrieve and extract information related to ship subcategories and their distinctive features from both internal and external knowledge sources. For instance, the Queen Elizabeth-class aircraft carrier is equipped with two ship islands: this type of feature can serve as crucial evidence in subsequent reasoning and decision-making processes and directly affects the system’s interpretation performance. In practical terms, knowledge retrieval involves two core technical challenges: efficient knowledge encoding and reliable knowledge retrieval. The former requires the model to embed various forms of data–such as structured data, unstructured text, or temporal information–into its internal parameters while ensuring that the information is well organized and meaningfully associated to avoid knowledge confusion. The latter demands that the model be able to rapidly locate and retrieve information highly relevant to the current task from a vast internal and external knowledge space by leveraging the complex associative relationships among knowledge elements.
III-A3 Decision-Making and Execution
Accurate decision-making and execution are the critical links driving the cyclic interactive interpretation process and bear the responsibility of translating cognitive reasoning from the global workspace into actual physical or digital operations. At this stage, the interpretation system must schedule matching interpretation tools from a capability library or issue specific control instructions to the observation system based on the task intent generated by reasoning. This process is not a simple unidirectional transmission of instructions but a complex mapping facing the dual challenges of the semantic gap and environmental dynamics. Specifically, this phase confronts the following two core challenges:
Tool Decision-Making
The primary challenge facing the interpretation system is bridging the semantic gap between natural language reasoning and rigid tool interfaces. Under the language-centered paradigm, LLM reasoning results typically exist in abstract natural language forms (e.g., “count the number of vessels in the port”), whereas existing remote sensing tools (such as object detection models, edge extraction algorithms, and change detection networks) rely on specific code invocation formats and parameter configurations. Enabling the system to accurately comprehend task intent and retrieve the most appropriate algorithm model from a massive, heterogeneous tool library is the core difficulty of tool decision-making. Furthermore, complex tasks often require a chained combination of multiple tools (e.g., performing “cloud removal” prior to “land-cover classification”), so the system must be able to plan a “chain of tools.” Therefore, designing a decision mechanism capable of precisely mapping unstructured language instructions to executable code or API calls is a key prerequisite for endowing the interpretation system with “operational agency”.
Action Execution
Upon determining “which tool to use” or “what observation to make,” designing an efficient execution mechanism to cope with the dynamically changing physical environment constitutes another major challenge. Remote sensing observation environments are typically characterized by high dynamism and are subject to uncontrollable factors such as cloud cover, illumination changes, and target movement, which can render preset perception strategies ineffective during execution. For instance, in UAV inspection or dynamic satellite scheduling, the system cannot rely solely on static task planning for information acquisition but must possess the ability to respond to environmental feedback in real time. This requirement implies that the interpretation system must establish an “input-feedback-correction” closed loop during action execution, dynamically adjusting sensor parameters (e.g., angle and resolution) or correcting flight paths based on the returned observation status. Consequently, constructing an adaptive execution mechanism that accommodates dynamic environmental constraints to ensure the robust acquisition of high-quality data serves as the physical foundation for realizing efficient reasoning and decision-making.